OpenStudy (anonymous):

NEED HELP!! Is (0,1) a solution to x+y=1?

OpenStudy (texaschic101):

(0,1) x = 0 and y = 1 x + y = 1 0 + 1 = 1 1 = 1 (correct) yes...it is a solution

OpenStudy (anonymous):

y=-x+1 1=-(0)+1 1= 1 yes

OpenStudy (anonymous):

yes it is a solution y=-x+1 1=-(0)+1 1=1

OpenStudy (anonymous):

just kidding